OpenStudy (anonymous):

solve the system of equations by using the elimination method. {4x^2+3y^2=12 {5x^2+6y^2=30

OpenStudy (campbell_st):

multiply the 1st equation by 2 and subtract the 2nd equation, this will eliminate y^2 so you have \[8x^2 + 6y^2 = 24\] \[-5x^2 - 6y^2 = -30\] ----------------------- \[3x^2 = -6\] this is interesting as you will have a complex solution. when you have x, substitute both values to find y.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

got it no solution